Suppose a firm’s profit is given by the equation p = –200 + 80Q – .2Q2....

  1. Suppose a firm’s profit is given by the equation p = –200 + 80Q – .2Q2. Which of the following is true?

a) The firm’s marginal profit is given by the equation: Mp = 80 – .2Q.

b) The firm’s profit-maximizing output is Q = 400.

c) The firm’s profit-maximizing output is Q = 200.

d) The firm’s marginal profit is given by the equation: Mp = 80 – 2Q.

e) The firm’s profit-maximizing output is Q = 800.

I know the answer is C; however I don't understand how to get there. Please include details/explanations and steps.
Answer #1

Profit = P = dP -2out 80 Q-0202 - 80 - 004 Q o= 80-004 Q 0.42 = 80 Q = 200/- pofit-maximizing outputplease thumbs up if it helped you.